OpenStudy (rissyroo118):

Which is an equation of the line containing the points (4, 10) and (6, 11) in standard form? –x + 2y = 16 2x – y = –2 x + 2y = 24 x – 2y = –16

OpenStudy (marissalovescats):

Well first you have to find your slop with the two points

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 10}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 11}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (rissyroo118):

okay so y - 10= m (x-4) ?